A dose of methylnaltrexone bromide is administered to patients weighing over 251 lbs at a level of 0.15 mg per kilogram of body mass. Each dose is administered through a subcutaneous injection every other day. If a patient weighs 264 lbs and the methylnaltrexone bromide solution has a concentration of 20 mg/ml, how many milliliters will be administered in a period of two weeks?
Answer #1

ANSWER.

Calculation of amount of drug administered in a period of 2 weeks.

                Ordered dose = 0.15 mg/kg.

            Weight of the patient = 264 lbs.

            Convert lbs value to kg for calculation.

                        1 kg = 2.205 lbs.

              So

                 Weight of the patient = 264/2.205

                                                   = 119.7 kg.

                                                   = 120 kg.

              Formula = Desired dose /Dose in hand * Quantity.

           Desired dose for a 120 kg patient(D) = 0.15 * 120

                                                                            = 18 mg.

                     Dose in hand (H) = 20 mg.

                       Quantity (Q) = 1 ml.

            As per formula

                  Amount (ml) of drug administered per day = D/H *Q

                                                                                      = 18/20 *1

                                                                                      = 0.9 ml.

           Amount (ml) of drug administered for a period

                                                                of 2 weeks = 0.9 * 14

                                                                              = 12.6 ml.